How much ammonia (NH3) could be melted by the addition of 3.5 kJ of heat, if the temperature of the liquid is at the melting point? (ΔHfus=5.66 kJ/mol)
1
Expert's answer
2016-04-25T12:18:05-0400
Q=ΔHfus·n
n=m/M
M (NH3)=17 g/mol
m (NH3)=(Q·M)/ΔHfus
m (NH3)=(3.5·17)/5.66=10.5 g